# Insights from leptohadronic modelling of the brightest blazar flare

### v1

@@ 4.2 Results of the “moving-blob” leptonic model
The failure to describe the flare in the framework of the “moving-blob” model implies that the observed emission during the flare originates in a quasi-stationary “standing feature” in the jet which does not move along with the jet plasma bulk motion and into which multiple blobs are injected as was modelled in Sect. 3.1. We discuss possible interpretations of the prevalence of the “standing feature” in Sect. 5.2.
4.3 Results of the leptohadronic model
@@
Figure 12: The light curve of 3C 454.3 based on the Fermi-LAT $\geq 100$ MeV photon flux (blue data points with error bars for detections if the source-detection test statistic $TS\geq 4$ and triangles for ULs otherwise, left ordinate axis) from the Fermi-LAT LCR (Abdollahi et al., 2023), and the corresponding IceCube daily muon neutrino yield for energies $E^{o}_{\nu}\geq 100$ TeV (grey curve, right ordinate axis). The Fermi-LAT flare threshold is defined as the 68% percentile of the distribution of the Fermi-LAT weekly-binned photon fluxes. The mean yearly IceCube muon neutrino rate for $E^{o}_{\nu}\geq 100$ TeV is shown in the legend along with the rates averaged over the periods of outbursts highlighted with grey and brown shaded regions. The IceCube muon neutrino alert IC120523A is shown as the dashed vertical curve.
@@
The plot with the leptohadronic “standing-feature” model fitting the observed SED for MJD 55517–55524 is shown in Fig. 11. In the leptohadronic model, all relevant processes responsible for the production of secondaries are considered: BH pair production, decay of particles produced in $p\gamma$ collisions, $\gamma\gamma$ absorption with the subsequent synchrotron and IC emission of the $e^{+}e^{-}$ pairs produced in these processes. Photon and all-flavour neutrino SEDs produced as a result of these processes are shown in Fig. 11 as colourful shades (see the legend) with their upper boundary corresponding to the UL proton energy density $u^{\prime\mathrm{UL}}=\rho_{p/e}u^{\prime}_{e}$ with $\rho_{p/e}=1.3\times 10^{2}$ and the lower boundary corresponding to the purely leptonic model with $u^{\prime}_{p}=0$. Proton and muon synchrotron and IC emissions are completely negligible. As in the case of TXS 0506+056 (e.g., Keivani et al., 2018), it is X-ray data which drive the constraints on $\rho_{p/e}$ where at higher values of $\rho_{p/e}$ the synchrotron emission from BH electrons starts overshooting the observed X-ray SED.
@@
The obtained proton energy density $u^{\prime\mathrm{UL}}_{p}=\rho_{p/e}u^{\prime}_{e}=1.1\times 10^{2}$ erg/cm3 corresponds to apparent jet power of $P^{\mathrm{UL}}_{p}=\pi c\beta\Gamma^{2}R^{\prime 2}_{b}u^{\prime\mathrm{UL}}_{p}=1.6\times 10^{48}$ erg/s while the Eddington luminosity is $L_{\mathrm{Edd}}=2.1\times 10^{47}$ erg/s. The fact that the apparent proton jet power exceeds the Eddington luminosity by an order of magnitude implies that these parameters cannot be at this level for a long time if we do not assume the possibility of super-Eddington accretion. Such high power is mainly attributed to a high value of the Lorentz factor $\Gamma=32$ during the flare and is lower during quieter periods of activity (see Appendix C). Alternatively, the value of the proton-to-electron-energy density can be lower than the UL $\rho_{p/e}=1.3\times 10^{2}$ or the SMBH mass is higher than we assumed. The derived value of $\rho_{p/e}$ is, however, two orders of magnitude lower than those obtained in some models for TXS 0506+056 (e.g., Cerruti et al., 2019) and of the same order of magnitude as those obtained by Keivani et al. (2018, model LMBB2b).
@@
[…]
by which we multiply the upper boundary of the neutrino SED presented in Fig. 11 assessed at the intermediate energy bins $E^{o}_{i}$ between $E^{o\,\mathrm{min}}_{\nu}$ and $E^{o\,\mathrm{max}}_{\nu}$ with the number of bins $N_{\mathrm{bins}}=30\times\lg(E^{o\,\mathrm{max}}_{\nu}/E^{o\,\mathrm{min}}_{\nu})$. $A_{\mathrm{eff}}(E^{o}_{\nu};\delta_{3C454.3})$ is the IceCube muon neutrino effective area at the declination of 3C 454.3 $\delta_{3C454.3}=16.15^{\circ}$ (IceCube Collaboration, 2021; IceCube Collaboration et al., 2021; Aartsen et al., 2020). The number $N_{\mathrm{max}}^{90\%\mathrm{\,CL}}=2.44$ (muon neutrinos per whole energy range from $E^{o\,\mathrm{min}}_{\nu}$ to $E^{o\,\mathrm{max}}_{\nu}$) is the UL at 90% confidence level from Feldman & Cousins (1998, table XII) in case of zero background and zero observed events. The time period considered is $\Delta T=\mathrm{1\,yr}$. The factor of $1/3$ accounts for neutrino oscillations in vacuum. Thus, the black solid curve with a downward arrow in Fig. 11 shows a one-year IceCube sensitivity [90% confidence UL] for the 3C 454.3 neutrino spectrum presented as the black shaded region in Fig. 11.
@@
4.4.1 On the IceCube neutrino alert IC120523A associated with 3C 454.3
@@
In IceCat-1 (Abbasi et al., 2023a), there is one muon neutrino event with estimated energy $E^{o}_{\nu}=168$ TeV and signalness (probability of the neutrino being astrophysical) $0.49$ positionally associated with 3C 454.3 albeit not contemporaneous with the studied flare. Taking the mean rate of $6\times 10^{-3}$ muon neutrinos per year from our estimations for 3C 454.3 and the duration of IceCube observations of $\approx 10$ years, we obtain the Poisson probability of detecting one or more muon neutrinos with energies $\geq 100$ TeV from 3C 454.3 to be $\left(1-\exp[-6\times 10^{-3}\times 10]\right)\approx 0.06$. The obtained value indicates that on a long-period average, our model agrees with the single positional association of a muon neutrino with 3C 454.3, especially in light of the possible Eddington bias for cosmic neutrino sources (Strotjohann et al., 2019). The fact that the neutrino event was on MJD 56070.6 during the global minimum of the $\gamma$-ray activity (see Fig. 12) may indicate (assuming the neutrino is not atmospheric) that the neutrino emission in blazars is lagging behind the leptonic electromagnetic emission (e.g., due to slowness of photopion processes, see Fig. 7). If we assume the protons continue to lose their energy under the same conditions as during MJD 55517–55524, then 99% of the neutrino fluence would be emitted in less than 450 days after the peak of the flare, while the alert came 551 days after the flare which makes the association of this neutrino with the flare improbable. However, if the Lorentz factor of the emitting region drops to $\Gamma\approx 10$ (typical for more quiescent states, see Appendix C), 99% of the neutrino fluence would be emitted in $\sim 1300$ days which covers well the neutrino-flare lag of 551 days. On the other hand, the expected neutrino yield from a source with a lower Lorentz factor is lower. If there is a time lag between neutrino and $\gamma$-ray emission, it may explain why 3C 454.3 was among the three brightest neutrino candidate sources in the search performed by Abbasi et al. (2023b) with $\gamma$-ray average flux weights (in agreement with our long-term average estimate) but was not among the best ten candidates in their search with monthly flux weights since they did not take into account a possible $\nu-\gamma$ time lag. The detailed investigation of the properties of delayed neutrino emission is beyond the scope of the paper.

### v2

@@ 4.2 Results of the “moving-blob” leptonic model
The failure to describe the flare in the framework of the “moving-blob” model implies that the observed emission during the flare originates in a quasi-stationary “standing feature” in the jet, which does not move along with the jet plasma bulk motion and into which multiple blobs are injected as was modelled in Sect. 3.1. We discuss possible interpretations of the prevalence of the “standing feature” in Sect. 7.2.
5 Description of the leptohadronic model
In this subsection, we obtain constraints on the maximum contribution of protons to the observed SED allowed by the data assuming the hadronic contribution in the model serves as a subdominant component of the model SED, in addition to the already obtained leptonic fit, similar to previous studies by, e.g., Keivani et al. (2018); Cerruti et al. (2019); Gao et al. (2019).
For the best fit of the “standing-feature” leptonic model averaged over MJD 55517–55524, we show the characteristic timescales in the blob frame in Fig. 12. To minimise the number of free parameters, we assume that the protons are co-accelerated with electrons in a process with an acceleration timescale
where $t^{\prime\mathrm{gyro}}_{p}=t^{\prime\mathrm{gyro}}_{e}$ is defined in Eq. (9). We assume for simplicity that $\xi_{\mathrm{acc}}$ is the same for both electrons and protons (see Appendix E for an investigation of other values of $\xi_{\mathrm{acc}}$ for protons). To estimate the value of $\xi_{\mathrm{acc}}$, we use the best-fit value of the maximum electron injection energy $E^{\prime\mathrm{max}}_{e}$ obtained from fitting the “standing-feature” leptonic model SED to the observed seven-day-averaged data (see Sect. 4.1) and find in Fig. 12 the interception of the vertical orange dashed-dotted line corresponding to $E^{\prime\mathrm{max}}_{e}$ (with the pale orange band representing its $1\sigma$ uncertainty) and the blue solid curve of the total electron energy-loss timescale, and superimpose the dependence of the gyroperiod on the energy (green dashed line) multiplied by $\xi_{\mathrm{acc}}=10^{7}$ (green solid line), which results in the self-consistent estimate of $E^{\prime\mathrm{max}}$ as the energy when the electron acceleration and energy-loss timescales become equal. This heuristically obtained approximate value of $\xi_{\mathrm{acc}}=10^{7}$ serves as the benchmark value for the leptohadronic model.
To determine the maximum proton energy, we assume that the acceleration started long before the flare [^18] Otherwise, the maximum proton energy would be determined by the interception of $t^{\prime\mathrm{acc}}_{p}$ with the simulation timescale $t^{\prime}_{\mathrm{sim}}$ and would be a few hundred GeV. and look at the interception of the green solid line of $t^{\prime\mathrm{acc}}_{p}(E^{\prime}_{p})$ with the proton escape timescale (grey solid line), as well as the proton energy-loss timescale $t^{\prime\mathrm{p\gamma}}$ due to photopion production (purple dotted line) having almost the same value at this energy (see Fig. 12):
$t^{\prime\mathrm{acc}}_{p}(E^{\prime{\mathrm{max}}}_{p})=\min\{t^{\prime p\gamma}_{p}(E^{\prime{\mathrm{max}}}_{p});t^{\prime\mathrm{esc}}_{p}\},$ (37)
where $t^{\prime\mathrm{esc}}_{p}=t^{\prime\mathrm{esc}}_{e}$ is defined in Eq. (11). This results in $E^{\prime{\mathrm{max}}}_{p}\approx 2\times 10^{14}\mathrm{\,eV}$ for MJD 55517–55524, where $E^{\prime{\mathrm{max}}}_{p}$ is taken as the energy of the exponential cutoff of the proton spectrum (shown as the orange dashed line in Fig. 12). We also check how this value varies from one day to another throughout the seven days: for a fixed value of $\xi_{\mathrm{acc}}=10^{7}$, the variation of $E^{\prime{\mathrm{max}}}_{p}$ is less than 3030%. The optimistic assumption that the proton acceleration started long before the flare is used, since we are focused on estimating the upper limit of the neutrino yield. The assumption implies that the observable neutrino energy should gradually rise in time from the beginning of the acceleration. However, due to the total expected number of observed neutrinos from one source $\ll 1$, this prediction is hardly testable even with the next generation neutrino telescopes. It is not common to expect that protons are accelerated long before electrons (because the acceleration timescale is proportional either to the particle energy or mass, Weidinger & Spanier, 2015, e.g.,), although it is not excluded since the protons, due to their significantly less efficient energy losses, are effectively decoupled from fast-cooling electrons and are not required to be continuously injected into the system as electrons. This assumption is not critical in our study, since it is used only in the heuristic way to find $E^{\prime{\mathrm{max}}}_{p}$, and a wide range of other possible values of the proton cutoff energy is considered in Appendix E. In an alternative assumption, when protons begin their acceleration at the beginning of the flare, $E^{\prime{\mathrm{max}}}_{p}$ could be defined as the energy at which the acceleration and flare (simulation) timescales become equal, in which case a different range of $\xi_{\mathrm{acc}}$ can be considered to cover the same range of $E^{\prime{\mathrm{max}}}_{p}$.
The minimal proton energy is fixed at $E^{\prime\mathrm{min}}_{p}=2\times 10^{9}$ eV, and their power-law spectral index fixed at $s_{p}=2.0$. In principle, the power-law spectral index $s_{p}$ of the proton distribution should follow the electron spectral index $s_{e}$, but, as can be seen in Fig. 8, $s_{e}$ varies noticeably from day to day, so we choose the standard value of $s_{p}=2$ for simplicity. Besides, in some acceleration scenarios the difference of 0.10.40.1-0.4 between $s_{e}$ and $s_{p}$ can be expected (e.g., Diesing & Caprioli, 2019). The uncertainty of the value of $\xi_{\mathrm{acc}}$ has a more significant impact on the neutrino SED than the uncertainty of $s_{p}$ [^19] E.g., Keivani et al. (2018) show how different values of $s_{p}$ affect the neutrino SED., because $\xi_{\mathrm{acc}}$ defines the peak energy of the neutrino SED as shown in Appendix E.
We choose the constant escape timescale with the value of $\eta_{p}=\eta_{e}\sim 5\times 10^{3}$ to avoid the dependence of the derived proton cutoff energy on a particular mechanism of the proton acceleration: in a case of a more efficient proton acceleration with $\xi_{\mathrm{acc}}<10^{7}$ the cutoff proton energy will be defined by the proton energy-loss timescale. We also verify that the most optimistic Bohm-limit diffusion escape timescale described by Eq. (10) and shown as the solid pink line in Fig. 12 lies somewhat higher than the proton energy-loss timescale, i.e. our approach of the heuristic determination of $E^{\prime{\mathrm{max}}}_{p}$ is not affected by a particular mechanism of the proton acceleration if $\xi_{\mathrm{acc}}\leq 10^{7}$. We explore how the observed electromagnetic and neutrino SEDs change if the proton acceleration is more efficient with $\xi_{\mathrm{acc}}<10^{7}$ in Appendix E.
drops to $BF=10^{-2}$ ($\approx 3\sigma$ Gaussian equivalent). We denote the corresponding value of the proton energy density $u^{\prime\mathrm{UL}}_{p}$ as the upper limit (UL). The obtained ratio between the UL proton energy density and injected electron energy density is (see Sect. 6)
The results obtained with the “standing-feature” leptohadronic model are presented in Sect. 6. Since the leptonic model within the “moving-blob” approach resulted in a poor description of the data (see Sect. 4.2), we do not develop its leptohadronic counterpart.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    The obtained with AM3 observable neutrino $\mathrm{SED}_{\nu}(E^{o}_{\nu})=E^{o\,2}_{\nu}dN^{o}_{\nu}/dE^{o}_{\nu}$ for each of the five considered periods is used to calculate the integral neutrino particle flux $\Phi^{o}_{\nu}$ for observable neutrino energies $E^{o}_{\nu}\geq 100$ GeV.
  5. 5.
    We obtained (see Appendix D) a good linear fit in the form of $\Phi^{o}_{\nu}=a\Phi^{o}_{\gamma}$, where $a$ is a constant, and $\Phi^{o}_{\gamma}$ is the integral model photon flux for observable $\gamma$-ray energies $E^{o}_{\gamma}\geq 100$ MeV.
  6. 6.
  7. 7.
  8. 8.
The obtained IceCube neutrino light curve and expected mean neutrino yield are presented in Sect. 6.1.
To estimate the neutrino flux observed by IceCube at energies $E^{o}_{\nu}\geq 100$ TeV, we sum up all signalnesses of all the IceCube alert tracks202020excluding those giving a signal in the IceTop detector from the public IceCat-1 catalogue (Abbasi et al., 2023a). This results in $\approx 12$ “pure” astrophysical muon neutrinos observed by IceCube at energies $E^{o}_{\nu}\geq 100$ TeV.
The estimated contribution of the Fermi-LAT FSRQs to the IceCube neutrino flux is presented in Sect. 6.2.
6 Results of the leptohadronic model
Figure 13: The photon and neutrino SED with model curves for the leptohadronic “standing-feature” model with a fit to the observed SED averaged over MJD 55517–5524 with proton acceleration parameter $\xi_{\mathrm{acc}}=10^{7}$. Shaded regions correspond to the spread of proton luminosities $0\leq u^{\prime}_{p}\leq u^{\prime\mathrm{UL}}_{p}$, where $u^{\prime\mathrm{UL}}_{p}$ is the UL corresponding to the Bayes factor with respect to the best-fit leptonic model $BF=10^{-2}$ ($\approx 3\sigma$ Gaussian equivalent). The crimson solid curve with a downward arrow shows the IceCube one-year sensitivity (see Eq. 41) for the UL of 2.44 muon neutrinos between $E^{o\,\mathrm{min}}_{\nu}=10^{14}$ eV and $E^{o\,\mathrm{max}}_{\nu}=10^{16}$ eV.
Figure 14: The inset of Fig. 13 focusing on the X-ray energy range. The solid orange curve with black dots shows the leptonic fit to the data, while the grey band shows the additional hadronic contribution due to the Bethe-Heitler synchrotron photons and the synchrotron and IC radiation by $e^{+}e^{-}$ pairs produced in $\gamma\gamma$-initiated cascades. The upper boundary of the grey band corresponds to the case of proton energy density tripled w.r.t. the one we considered as the upper limit: $u^{\prime}_{p}=3\times u^{\prime\mathrm{UL}}_{p}$.
@@ 6 Results of the leptohadronic model
Figure 15: The light curve of 3C 454.3 based on the Fermi-LAT $\geq 100$ MeV photon flux (blue data points with error bars for detections if the source-detection test statistic $TS\geq 4$ and triangles for ULs otherwise, left ordinate axis) from the Fermi-LAT LCR (Abdollahi et al., 2023), and the corresponding IceCube daily muon neutrino yield for energies $E^{o}_{\nu}\geq 100$ TeV (grey curve, right ordinate axis). The Fermi-LAT flare threshold is defined as the 68% percentile of the distribution of the Fermi-LAT weekly-binned photon fluxes. The mean yearly IceCube muon neutrino rate for $E^{o}_{\nu}\geq 100$ TeV is shown in the legend along with the rates averaged over the periods of outbursts highlighted with grey and brown shaded regions. The IceCube muon neutrino alert IC120523B is shown as the dashed vertical curve.
@@
The plot with the leptohadronic “standing-feature” model fitting the observed SED for MJD 55517–55524 is shown in Fig. 13. In the leptohadronic model, all relevant processes responsible for the production of secondaries are considered: BH pair production, decay of particles produced in $p\gamma$ collisions, $\gamma\gamma$ absorption with the subsequent synchrotron and IC emission of the $e^{+}e^{-}$ pairs produced in these processes. Photon and all-flavour neutrino SEDs produced as a result of these processes are shown in Fig. 13 as colourful shades (see the legend) with their upper boundary corresponding to the UL proton energy density $u^{\prime\mathrm{UL}}_{p}=\rho_{p/e}u^{\prime}_{e}$ with $\rho_{p/e}=1.3\times 10^{2}$ and the lower boundary corresponding to the purely leptonic model with $u^{\prime}_{p}=0$. Proton and muon synchrotron and IC emissions are completely negligible. As in the case of TXS 0506+056 (e.g., Keivani et al., 2018), it is X-ray data which drive the constraints on $\rho_{p/e}$ where at higher values of $\rho_{p/e}$ the synchrotron emission from BH electrons as well as the synchrotron and IC radiation by $e^{+}e^{-}$ pairs produced in $\gamma\gamma$-initiated cascades start overshooting the observed X-ray SED. In Fig. 13, however, this cascade contribution to the X-ray flux seems to be subdominant, and it appears to the naked eye that an even higher $u^{\prime\mathrm{UL}}_{p}$ would still be allowed by the data. Similar to how it has been done by Keivani et al. (2018), we demonstrate in the zoomed-in X-ray area of the SED plot, how the model SED looks like in case $u^{\prime}_{p}=3\times u^{\prime\mathrm{UL}}_{p}$ as shown in Fig. 14. As seen in the X-ray inset, in the case when the proton energy density is three times greater than the one we consider as the upper limit $u^{\prime\mathrm{UL}}_{p}$, the model SED exceeds $1\sigma$ error bars in $12$ out of $14$ data points, clearly over-predicting the X-ray flux.
@@
The obtained proton energy density $u^{\prime\mathrm{UL}}_{p}=\rho_{p/e}u^{\prime}_{e}=1.1\times 10^{2}$ erg/cm3 corresponds to apparent jet power of $P^{\mathrm{UL}}_{p}=\pi c\beta\Gamma^{2}R^{\prime 2}_{b}u^{\prime\mathrm{UL}}_{p}=1.4\times 10^{48}$ erg/s while the Eddington luminosity is $L_{\mathrm{Edd}}=2.1\times 10^{47}$ erg/s. The fact that the apparent proton jet power exceeds the Eddington luminosity by an order of magnitude implies that these parameters cannot be at this level for a long time if we do not assume the possibility of super-Eddington accretion. Such high power is mainly attributed to a high value of the Lorentz factor $\Gamma=35$ during the flare and is lower during quieter periods of activity (see Appendix C). Alternatively, the value of the proton-to-electron-energy density can be lower than the UL $\rho_{p/e}=1.3\times 10^{2}$ or the SMBH mass is higher than we assumed. The derived value of $\rho_{p/e}$ is, however, two orders of magnitude lower than those obtained in some models for TXS 0506+056 (e.g., Cerruti et al., 2019) and of the same order of magnitude as those obtained by Keivani et al. (2018, model LMBB2b).
@@
$m_{i}=\frac{N_{\mathrm{max}}^{90\%\mathrm{\,CL}}}{N_{\mathrm{bins}}}\left[\int_{E^{o}_{i-1}}^{E^{o}_{i}}\frac{1}{3}\frac{dN_{\nu}^{o}}{dE^{o}_{\nu}}A_{\mathrm{eff}}(E^{o}_{\nu};\delta_{\mathrm{obj}})dE^{o}_{\nu}\times{\Delta T}\right]^{-1}$ (41)
by which we multiply the upper boundary of the neutrino SED presented in Fig. 13 assessed at the intermediate energy bins $E^{o}_{i}$ between $E^{o\,\mathrm{min}}_{\nu}$ and $E^{o\,\mathrm{max}}_{\nu}$ with the number of bins $N_{\mathrm{bins}}=30\times\lg(E^{o\,\mathrm{max}}_{\nu}/E^{o\,\mathrm{min}}_{\nu})$. $A_{\mathrm{eff}}(E^{o}_{\nu};\delta_{\mathrm{obj}})$ is the IceCube muon neutrino effective area at the declination of 3C 454.3 $\delta_{\mathrm{obj}}=16.15^{\circ}$ (IceCube Collaboration, 2021; IceCube Collaboration et al., 2021; Aartsen et al., 2020). The number $N_{\mathrm{max}}^{90\%\mathrm{\,CL}}=2.44$ (muon neutrinos per whole energy range from $E^{o\,\mathrm{min}}_{\nu}$ to $E^{o\,\mathrm{max}}_{\nu}$) is the UL at 90% confidence level from Feldman & Cousins (1998, table XII) in case of zero background and zero observed events. The time period considered is $\Delta T=\mathrm{1\,yr}$. The factor of $1/3$ accounts for neutrino oscillations in vacuum. Thus, the crimson solid curve with a downward arrow in Fig. 13 shows a one-year IceCube sensitivity [90% confidence UL] for the 3C 454.3 neutrino spectrum presented as the crimson shaded region in Fig. 13.
@@
Capel et al. (2022) found that, to be consistent with a single coincident observation of a neutrino alert and flaring $\gamma$-ray blazar (i.e. the association of the flare in TXS 0506+056 with the neutrino alert IC170922A, IceCube Collaboration et al., 2018), for FSRQs, the ratio $Y_{\nu\gamma}$ (see their eq. 18 and fig. 8) of the neutrino energy flux in the range $10$ TeV–$100$ PeV over the $\gamma$-ray energy flux in the range 11100100 GeV must be $10^{-5}\lesssim Y_{\nu\gamma}\lesssim 10^{-2}$ if neutrinos are produced in both quite and flaring states of FSRQs, or $10^{-4}\lesssim Y_{\nu\gamma}\lesssim 10^{0}$ if neutrinos are produced only during flares. In our fiducial leptohadronic model for MJD 55517–55524, we find $Y_{\nu\gamma}\approx 8\times 10^{-4}$, which is in good agreement with both constraints by Capel et al. (2022).
6.1.1 On the IceCube neutrino alert IC120523B associated with 3C 454.3
@@
In IceCat-1 (Abbasi et al., 2023a), there is one muon neutrino event with estimated energy $E^{o}_{\nu}=168$ TeV and signalness (probability of the neutrino being astrophysical) $0.49$ positionally associated with 3C 454.3 albeit not contemporaneous with the studied flare [^21] In the original version of the catalogue presented in the article by Abbasi et al. (2023a) this event due to a typo was called IC120523A, having the same name as another alert. However, in the updated version of IceCat-1 (IceCube Collaboration, 2023), the alert is called IC120523B.. Taking the mean rate of $6\times 10^{-3}$ muon neutrinos per year from our estimations for 3C 454.3 and the duration of IceCube observations of $\approx 10$ years, we obtain the Poisson probability of detecting one or more muon neutrinos with energies $\geq 100$ TeV from 3C 454.3 to be $\left(1-\exp[-6\times 10^{-3}\times 10]\right)\approx 0.06$. The obtained value indicates that on a long-period average, our model agrees with the single positional association of a muon neutrino with 3C 454.3, especially in light of the possible Eddington bias for cosmic neutrino sources (Strotjohann et al., 2019). The fact that the neutrino event was on MJD 56070.6 during the global minimum of the $\gamma$-ray activity (see Fig. 15) may indicate (assuming the neutrino is not atmospheric) that the neutrino emission in blazars is lagging behind the leptonic electromagnetic emission (e.g., due to slowness of photopion processes, see Fig. 12). If we assume the protons continue to lose their energy under the same conditions as during MJD 55517–55524, then 99% of the neutrino fluence would be emitted in less than 450 days after the peak of the flare, while the alert came 551 days after the flare, which makes the association of this neutrino with the flare improbable. However, if the Lorentz factor of the emitting region drops to $\Gamma\approx 10$ (typical for more quiescent states, see Appendix C), 99% of the neutrino fluence would be emitted in $\sim 1300$ days, which covers well the neutrino-flare lag of 551 days. On the other hand, the expected neutrino yield from a source with a lower Lorentz factor is lower. If there is a time lag between neutrino and $\gamma$-ray emission, it may explain why 3C 454.3 was among the three brightest neutrino candidate sources in the search performed by Abbasi et al. (2023b) with $\gamma$-ray average flux weights (in agreement with our long-term average estimate) but was not among the best ten candidates in their search with monthly flux weights since they did not take into account a possible $\nu-\gamma$ time lag. The detailed investigation of the properties of delayed neutrino emission is beyond the scope of the paper.